OpenStudy (anonymous):

A line has a slope of −5 and contains the point (−2, 4). Which equations represent the line? Choose ALL answers that are correct. ▢A.−5x + y = −6 ▢B.5x + y = −6 ▢C.(y – 4) = −5(x + 2) ▢D.(y + 4) = −5(x − 2) I think the answers are A and C but I'm not sure .. Check ? an hour ago

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

y=mx+b, slope m = -5, so y = -5x+b now we can supstiture x= -2 and y = 4 \[4 = -5 \times -2 +b\] hence \[b = -6\] so the line can be expressed as y = -5x -6 this is equal to -5x + y = -6 (by adding -5x to both sides) -> A this is not the same as B if we expand expression C we get \((y – 4) = −5(x + 2)\), or \((y – 4) = −5x -10)\), that is \(y = −5x -6\), so C is also equivalent D \((y + 4) = −5(x − 2)\) is equivalent to \(y + 4 = −5x + 10\), that is \(y = −5x + 6\) So D is not valid You were correct

OpenStudy (anonymous):

@kewlkatcutie

OpenStudy (anonymous):

Thank you so sorry for the late response <33

OpenStudy (anonymous):

It's ok, it is not that late either